# OplixNet: Towards Area-Efficient Optical Split-Complex Networks with Real-to-Complex Data Assignment and Knowledge Distillation

### 03 Dec 2023 v1

@@ I Introduction
Artificial neural networks (ANN) have garnered significant achievements across numerous domains, including computer vision [1] and natural language processing [2]. Recently, the development of emerging large language models, such as ChatGPT, caused a higher demand for neuromorphic computing devices[3]. Besides CPUs, GPUs, TPUs [4] and other ASICs, some emerging devices, such as superconducting circuits [5], RRAM [6, 7, 8], and optical devices [9] are also proposed to further boost computing speed, reduce energy consumption and save device area of ANN computations.
Optical neural network (ONN) [9, 10, 11, 12, 13] is a new neuromorphic computing solution based on optical components. The ONN structure proposed in [9] uses arrays of Mach-Zehnder Interferometers (MZIs) to perform the matrix-vector multiplications (MVMs). The weight matrices obtained from software-trained networks are mapped to the phases of the MZIs via singular value decomposition (SVD) and unitary-to-interferometer mapping [14]. The amplitudes and phases of the light signal are changed while going through an MZI array to implement multiplications and additions. These computations can be conducted in an extremely low latency with an optical detection rate over 100 GHz [15].
@@ Acknowledgement
This work is supported by TUM International Graduate School of Science and Engineering (IGSSE), Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 497488621, NSFC (Grant No. 62034007, 62141404) and SGC Cooperation Project (Grant No. M-0612).

### 15 Dec 2023 v2

@@ I Introduction
Artificial neural networks (ANN) have garnered significant achievements across numerous domains, including computer vision [1] and natural language processing [2]. Recently, the development of emerging large language models, such as ChatGPT, caused a higher demand for neuromorphic computing devices. Besides CPUs, GPUs, TPUs [3] and other ASICs, some emerging devices, such as superconducting circuits [4], RRAM [5, 6, 7, 8, 9], and optical devices [10, 11, 12, 13] are also proposed to further boost computing speed, reduce energy consumption and save device area of ANN computations.
Optical neural network (ONN) is a new neuromorphic computing solution based on optical components. The ONN structure proposed in [10] uses arrays of Mach-Zehnder Interferometers (MZIs) to perform the matrix-vector multiplications (MVMs). The weight matrices obtained from software-trained networks are mapped to the phases of the MZIs via singular value decomposition (SVD) and unitary-to-interferometer mapping [14]. The amplitudes and phases of the light signal are changed while going through an MZI array to implement multiplications and additions. These computations can be conducted in an extremely low latency with an optical detection rate over 100 GHz [15].
@@ Acknowledgement
This work is supported by TUM International Graduate School of Science and Engineering (IGSSE), Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 497488621, NSFC (Grant No. 62034007, 62141404), the Major Program of Zhejiang Provincial NSF D24F040002, and SGC Cooperation Project (Grant No. M-0612).